OpenStudy (anonymous):

Eve needs to purchase a fish tank with a volume of 3 cubic feet for her ruby discus. the tank must be 24 inches long and 18 inches wide. given these dimensions, what should be the height, in inches, of the tank? a. 14 inches b. 12 inches c. 39 inches d. 2 inches

OpenStudy (anonymous):

First convert feet to inches Then plug in and solve V= L * W * H

OpenStudy (anonymous):

27 = (1.5) (2) x all in feet x = 9

OpenStudy (anonymous):

Oops miss read 3= (1.5) (2) x all in feet x = 1 feet = 12 inch

OpenStudy (anonymous):

1ft=12in 1ft^3=12^3in^3=1728 in^3 1728=24*18*H H=1728/(24*18) H=4

OpenStudy (anonymous):

1 cubic feet is equal to 1728 cubic inches.. so 3*1728=24*18*h h=12 inches

OpenStudy (anonymous):

hahaha, my bad. I forgot to multiply by 3. LeopoldBr is right

OpenStudy (anonymous):

K thanks everybody :)